• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kramer

    Log entry time 00:01:22 on June21,2002

    Entry number 86529

    keyword=Shift Summary

    1600 take over, stable running
    1626 start run 1697 - efficiency run
    1633 start run 1698 - production run
    1648 end run 1698 for ep energy measurement
    1650 start ep energy measurements / change kinematics to Q2 g50
    1831 ep measurements done, going to optics target and beam
    1839 beam set, start optics run
    1840 start run 1699 - optics run - wrong prescale
    1847 start run 1700 - optics with correct prescale
    1854 move to d2 tgt/raster back on
    1910 start run 1701- first production run Q2 g50
    2000 beam lost, end run, coda crashes at run end, restart coda
    2026 start run 1702
    2122 start run 1703
    2205 daq crash ends run 1703, restart coda
    2224 start run 1704
    2327 start run 1705
    2332 daq problems end run 1705, all rocs not communicating
    2334 start run 1706